# On The Implicit Large Eddy Simulation of Turbomachinery Flows Using The Flux Reconstruction Method

### v1

@@ 4.3.1 Effect of DOFs
For a typical second-order finite volume solver, DOFs in the computation are determined by the grid size. However, for a spectral element solver (i.e. FR), both the grid size ($h$) and the polynomial order ($p$) can be adjusted to increase the DOFs in the computation. It is not computationally feasible to perform a sensitivity study of both the grid size (i.e. in streamwise, wall normal, and spanwise directions) and the polynomial order altogether for the LES of an industrial case. In this work, a segmented approach is used. A series of 2D calculations have been conducted in the first place to determine a desirable grid spacing in the streamwise ($x^{+}$) and the normal direction ($y^{+}$) on the blade surface for a $p3$ calculation. For simplicity, this 2D study is not shown here. The resulting 2D mesh is then used to perform the sensitivity study in 3D. This study is summarized in Table 6. Mean flow quantities that are relevant to engineering designs, such as skin friction, the isentropic Mach number ($\text{M}_{\text{is}}$) and HTC are used as criteria to demonstrate the effect of increasing the polynomial order $p$ and reducing the grid size $h$ on the solution, respectively.
@@
The HLLC Riemann solver is used in this sensitivity study and the choice of $s_{0}$ for LMF is guided by Table 2(b). $p2$ and $p3$ computations are performed on a mesh with 10 hexahedral layers in the spanwise direction. This is then increased to 15 and 20 layers. Figure 10 shows the isosurface of the instantaneous flow Q criterion of setup 4, and the isosurface is colored by the axial velocity. The flow is laminar on the pressure side but transitional on the suction side and then becomes fully turbulent towards the trailing edge. Such a flow structure is representative for all the setups. Figure 11 shows the representative $y^{+}$ on the blade for Setups 2,3 and 4. It can be seen that the averaged $y^{+}$ on the blade surface for $p3$ is approximately 0.75. Regarding $x^{+}$, which represents the grid spacing in the streamwise, they are approximately 30 for Setups 1-4. With respect to $z^{+}$, they are on average 90, 60 and 45 on the blade surface for Setups 1&\&&2, 3 and 4, respectively.
@@ 4.3.2 Effect of Riemann Solvers and Filters
The effect of the different setups in Table 7 on instantaneous flow can be qualitatively demonstrated by Fig 15. The figure shows the magnitude of the density gradient in the middle span section. All setups capture the dominant flow features of the unsteady flow, for example, the suction side transition, vortex shedding from the trailing edge, and the resulting pressure waves radiating from the trailing edge. HLLC-LMF shows a resolution of the pressure waves comparable to Roe-LMF. With a more relaxed smoothness criterion, HLLC-LMF-2 further improves the resolution of the pressure waves compared to HLLC-LMF and Roe-LMF.
@@
With regard to EF, it successfully stabilizes the solution without any tuning parameters, but shows a less sharp resolution of the pressure waves compared to HLLC-LMF-2. This observation is consistent with the results reported by Gao et al. [54], who reported that EF could introduce slightly more dissipation into the computation. This observation can be confirmed by Fig. 16. The PSD of the fluctuation of pressure is shown at a point that is located downstream of the trailing edge and is in the mid-span plane. Its coordinate on the blade-to-blade section is $(0.03722,-0.05482)$. The frequency is represented in terms of the Strouhal number. The characteristic length $L_{char}$ is the radius of the trailing edge, and the characteristic velocity $u_{char}$ is the magnitude of the velocity at the exit boundary. Based on the selected time step, the cut-off Strouhal number is $St_{\text{cut-off}}\approx 1000$. For clarity, only the comparison of HLLC-LMF-2 and HLLC-EF is shown, and the PSD of other setups with LMF in Table 7 is found to be very similar to HLLC-LMF-2. From the figure, it can be seen that EF shows a faster decay of PSD at higher frequencies than HLLC-LMF-2 and this indicates that there is more dissipation in the calculation of HLLC-EF than that of HLLC-LMF-2. In addition, it is interesting to observe that the PSD of EF is less noisy than HLLC-LMF-2 at higher frequencies. Further investigation is required to understand this scenario.
@@ 4.4 Remarks on the filter and smoothness criterion
Table 2(b) shows the range of $s_{0}$ in which LMF could potentially be used to stabilize the computation and has been used as a guide to choose the values of $s_{0}$ in the T106c and LS89 turbine cases. Here, we perform a posteriori analysis of the smooth criterion using the values of $s_{e}$ for the T106c and LS89 turbine cases. To the best knowledge of the author, there have been no previous work showing the variation of flow smoothness $s_{e}$ for turbomachinery flows. Therefore it is useful to conduct such a analysis as it can provide a rationale for the range of $s_{0}$ in Table 2(b) and guide the selection of a suitable $s_{0}$.
@@ 4.4.2 The LS89 Turbine
Figure 19 shows the contours of the smooth indicator $s_{e}$ for the setups in Table 7 that uses LMF. The plots in the bottom row highlight the cells that are subject to filtering. From the plot, it can be seen that the smooth indicator can identify the regions where the flow is smooth and also the regions where the flows have a large spatial gradient. Compared to the T106c case, more cells are filtered due to the more complicated flow features in this case. However, only a tiny proportion of elements are filtered. This confirms that LMF introduces a very low dissipation. This demonstrates the effectiveness of the LMF in stabilizing the solution while leaving the smooth-flow region untouched. Furthermore, Fig. 19 shows that the value of $s_{e}$ in the wake region and also in the transitional flow region is approximately in the range of $[-5:-4]$. This is consistent with the range suggested in Table 2(b).
@@
Regarding EF, $\alpha$ in the second order modal filter used in EF is shown in Figure 21. In EF, $\alpha$ is adaptively computed based on physical constraints (e.g., the minimum entropy principle and the positivity-preserving condition). When $\alpha=0$, no filtering is applied, it can be seen from Fig. 21 that the physical constraint in EF is capable of detecting the smooth flow region. When $\alpha>0$, the second modal filter is applied to the element. Compared with Fig. 19 and Fig. 21, it can be seen that both EF and LMF target similar flow regions, but EF filters more elements than LMF, although the value of $\alpha$ is small for most elements that are subjected to filtering.
@@ 4.4.3 Remarks on LMF and EF
As a comparison between LMF and EF, in terms of computational cost, EF requires an optimization process to obtain a suitable $\alpha$ in Equation 20 for each cell. For LMF, only the smoothness indicator (see Equation 19) needs to be computed, which is very efficient to evaluate. Therefore, LMF is computationally more efficient than EF. For the LS89 turbine case, each EF iteration is found to be roughly $15\%$ more expensive than that of LMF. In addition, since LMF is also easier to implement than EF, LMF can be a simple yet effective alternative to EF. On the other hand, EF is not completely parameter free. It still has a tuning parameter ${\epsilon}_{s}$ (see Equation 21) and this parameter controls how strict the minimum entropy principle is preserved. More research is still required to understand how this parameter can be tuned to strike a balance between accuracy and stability. However, it should be noted that LMF and EF still follow the same strategy, which is to localize the modal filtering process to strike a balance between accuracy and robustness.
@@ 5 Conclusions and Future Work
An FR solver ATHOS has been developed and validated to simulate turbomachinery flows. Its performance is demonstrated in two industrially representative turbomachinery cases, and the results show excellent agreement with the experimental data. LMF has been developed to stabilize the solution, and the TGV case shows that LMF does not introduce a noticeable amount of numerical dissipation when the flow is smooth. The choice of Riemann solvers (i.e. HLLC or Roe) can have an impact on the non-linear stability of the solution; HLLC is found to be more robust than Roe. However, if stable results can be obtained, only marginal differences are observed between the results of HLLC and Roe, and the quality of the solution is dominated more by the smoothness criterion $s_{0}$ than by the choice of the Riemann solver. A more relaxed spanwise grid spacing can be used to predict the HTC of the turbine blades. As the industry is more interested in mean flow quantities, this finding could lead to a reduction in the computational cost of industrial ILES simulations.
@@
Future work includes exploring the possibility of using other flow variables to calculate the smoothness indicator. In order to further enhance the robustness, a more advanced approach could be used to compute the convective flux (such as an entropy stable scheme). Besides, the connection of EF or LMF to SVV within the FR framework could be explored, this could gain more insight into the behavior of EF and LMF and leads to potential future improvement. As was pointed out by Edoh et al. [43], solution-filtering approach can potentially lead to time inconsistency, this will also be investigated in the future work.

### v2

@@ 4.3.1 Effect of DOFs
For a typical second-order finite-volume solver, DOFs in the computation are determined by the grid size. However, for a spectral element solver (i.e. FR), both the grid size ($h$) and the polynomial order ($p$) can be adjusted to increase the DOFs in the computation. It is not computationally feasible to perform a sensitivity study of both the grid size (i.e. in streamwise, wall normal, and spanwise directions) and the polynomial order altogether for the LES of an industrial case. In this work, a segmented approach is used. In the first place, a series of 2D calculations have been performed to determine a desirable grid spacing in the streamwise ($x^{+}$) and the normal direction ($y^{+}$) on the blade surface for a $p3$ calculation. For simplicity, this 2D study is not shown here. The resulting 2D mesh is then used to perform the sensitivity study in 3D. This study is summarized in Table 6. Mean flow quantities that are relevant to engineering designs, such as skin friction, the isentropic Mach number ($\text{M}_{\text{is}}$) and HTC are used as criteria to demonstrate the effect of increasing the polynomial order $p$ and reducing the grid size $h$ on the solution, respectively.
@@
The HLLC Riemann solver is used in this sensitivity study and the choice of $s_{0}$ for LMF is guided by Table 2(b). $p2$ and $p3$ computations are performed on a mesh with 10 hexahedral layers in the spanwise direction. This is then increased to 15 and 20 layers. Figure 10 shows the isosurface of the instantaneous flow Q criterion of Setup 4, and the isosurface is colored by the axial velocity. The flow is laminar on the pressure side but transitional on the suction side and then becomes fully turbulent towards the trailing edge. This flow structure is representative for all the setups. Figure 11 shows the representative $y^{+}$ on the blade for Setups 2,3 and 4. It can be seen that the averaged $y^{+}$ on the blade surface for $p3$ is approximately 0.75. Regarding $x^{+}$, which represents the grid spacing in the streamwise, they are approximately 30 for Setups 1-4. With respect to $z^{+}$, they are on average 90, 60 and 45 on the blade surface for Setups 1&\&&2, 3 and 4, respectively.
@@ 4.3.2 Effect of Riemann Solvers and Filters
The effect of the different setups in Table 7 on instantaneous flow can be qualitatively demonstrated by Fig 15. The figure shows the magnitude of the density gradient in the middle-span section. All setups capture the dominant flow features of the unsteady flow, for example, the suction side transition, vortex shedding from the trailing edge, and the resulting pressure waves radiating from the trailing edge. HLLC-LMF shows a resolution of the pressure waves comparable to Roe-LMF. With a more relaxed smoothness criterion, HLLC-LMF-2 further improves the resolution of the pressure waves compared to HLLC-LMF and Roe-LMF.
@@
With regard to EF, it successfully stabilizes the solution without any tuning parameters but shows a less sharp resolution of the pressure waves compared to HLLC-LMF-2. This observation is consistent with the results of Gao et al. [54], who reported that EF could introduce slightly more dissipation into the computation. In order to demonstrate this, the PSD of pressure is monitored for a point, which is located downstream of the trailing edge on the mid-span plane. Its coordinate on the blade-to-blade section is $(0.03722,-0.05482)$. The PSD for this coordinate was also computed by Jia et al. [8] and therefore their data is included. This comparison is shown in Fig. 16. The frequency is represented in terms of the Strouhal number. The characteristic length $L_{char}$ is the radius of the trailing edge and the characteristic velocity $u_{char}$ is the magnitude of the velocity at the exit boundary. Based on the selected time step, the cutoff Strouhal number is $St_{\text{cut-off}}\approx 200$. For clarity, only the comparison of HLLC-LMF-2 and HLLC-EF is shown, and the PSD of other setups with LMF in Table 7 is found to be very similar to HLLC-LMF-2. The results of Jia are computed by the code hpMusic, which implements the CPR method. ILES with implicit time stepping is used, and the cut-off Strouhal number is around 10.
From Jia’s CPR solutions, it can be seen that increasing the spatial accuracy from p2 to p3 shows a better agreement with the -5/3 rule on the inertial subrange. Jia’s data also show that more numerical dissipation leads to a faster decay of PSD, especially at higher frequencies. When comparing the CPR p3 solution with HLLC-LMF-2, both data agree well on the inertial subrange, but the CPR p3 solution shows a faster decay of PSD at higher frequencies. This indicates that HLLC-LMF-2 is less dissipative than Jia’s CPR p3 solution. Jia also compared the impact of explicit and implicit time stepping on PSD, the comparison showed that the choice of time-stepping did not have a noticeable impact on PSD. Therefore, the reason why Jia’s CPR p3 solution is more dissipative than the HLLC-LMF-2 result could be caused by the limiter used in hpMusic to stabilize the computation. With respect to EF, it is interesting to observe that its PSD is similar to that of Jia’s CPR p3 data. This shows that EF does introduce a slightly higher dissipation at higher frequencies than HLLC-LMF-2, but this does not seem to have a big impact on the inertial subrange.
@@ 4.4 Remarks on the filter and smoothness criterion
Table 2(b) shows the range of $s_{0}$ in which LMF could potentially be used to stabilize the computation and has been used as a guide to choose the values of $s_{0}$ in the T106c and LS89 turbine cases. Here, we perform a posteriori analysis of the smooth criterion using the values of $s_{e}$ for the T106c and LS89 turbine cases. To the best knowledge of the author, there has been no previous work showing the variation of flow smoothness $s_{e}$ for turbomachinery flows. Therefore, it is useful to perform such an analysis as it can provide a rationale for the range of $s_{0}$ in Table 2(b) and guide the selection of a suitable $s_{0}$.
@@ 4.4.2 The LS89 Turbine
Figure 19 shows the contours of the smooth indicator $s_{e}$ for the setups in Table 7 that use LMF. The plots in the bottom row highlight the cells that are subject to filtering. From the plot, it can be seen that the smooth indicator can identify the regions where the flow is smooth and also the regions where the flows have a large spatial gradient. Compared to the T106c case, more cells are filtered due to the more complicated flow features in this case. However, only a tiny proportion of elements are filtered. This confirms that LMF introduces a very low dissipation. This demonstrates the effectiveness of the LMF in stabilizing the solution while leaving the smooth-flow region untouched. Furthermore, Fig. 19 shows that the value of $s_{e}$ in the wake region and also in the transitional flow region is approximately in the range of $[-5:-4]$. This is consistent with the range suggested in Table 2(b).
@@
Regarding EF, $\alpha$ in the second-order modal filter used in EF is shown in Figure 21. In EF, $\alpha$ is adaptively computed based on physical constraints (e.g., the minimum entropy principle and the positivity-preserving condition). When $\alpha=0$, no filtering is applied, it can be seen from Fig. 21 that the physical constraint in EF is capable of detecting the smooth flow region. When $\alpha>0$, the second modal filter is applied to the element. Compared with Fig. 19 and Fig. 21, it can be seen that both EF and LMF target similar flow regions, but EF filters more elements than LMF, although the value of $\alpha$ is small for most elements that are subjected to filtering.
@@ 4.4.3 Remarks on LMF and EF
As a comparison between LMF and EF, in terms of computational cost, EF requires an optimization process to obtain a suitable $\alpha$ in Equation 20 for each cell. For LMF, only the smoothness indicator (see Equation 19) needs to be computed, which is very efficient to evaluate. Therefore, LMF is computationally more efficient than EF. For the LS89 turbine case, each EF iteration is found to be roughly $15\%$ more expensive than that of LMF. In addition, since LMF is also easier to implement than EF, LMF can be a simple yet effective alternative to EF. On the other hand, EF is not completely parameter-free. It still has a tuning parameter ${\epsilon}_{s}$ (see Equation 21) and this parameter controls how strict the minimum entropy principle is preserved. More research is still required to understand how this parameter can be tuned to strike a balance between accuracy and stability. However, it should be noted that LMF and EF still follow the same strategy, which is to localize the modal filtering process to strike a balance between accuracy and robustness.
@@ 5 Conclusions and Future Work
An FR solver ATHOS has been developed and validated to simulate turbomachinery flows. Its performance is demonstrated in two industrially representative turbomachinery cases, and the results show excellent agreement with the experimental data. LMF has been developed to stabilize the solution, and the TGV case shows that LMF does not introduce a noticeable amount of numerical dissipation when the flow is smooth. A more relaxed spanwise grid spacing can be used to predict the HTC of the turbine blades. As the industry is more interested in mean flow quantities, this finding could lead to a reduction in the computational cost of industrial ILES simulations.
The choice of Riemann solvers (i.e. HLLC or Roe) can have an impact on the non-linear stability of the solution; HLLC is found to be more robust than Roe. However, if stable results can be obtained, with respect to the mean flow only marginal differences are observed between the results of HLLC and Roe for the mean flow, and the predicted mean flow solution is dominated more by the smoothness criterion $s_{0}$ than by the choice of the Riemann solver; with respect to the instataneous flow, the PSD of pressure of the T106c turbine shows that the results from HLLC and Roe agree with each other in the inertial subrange but the PSD from the Roe solver produces a faster decay of PSD at higher frequencies. For the LS89 turbine case, all p3 solutions agree well with each other in the inertial subrange, a more dissipative stabilization approach, such as EF or the CPR limiter in Jia et al. [8] predicts a faster decay of PSD.
@@
Future work includes exploring the possibility of using other flow variables to calculate the smoothness indicator. In order to further enhance the robustness, a more advanced approach could be used to compute the convective flux (such as an entropy stable scheme). Besides, the connection of EF or LMF to SVV within the FR framework could be explored, this could gain more insight into the behavior of EF and LMF and leads to potential future improvement. As was pointed out by Edoh et al. [43], solution-filtering approach can potentially lead to time inconsistency; this will also be investigated in future work.